What is the slope of the line containing the points: (-4, -3), (2, 6), and (6, 12)
Answer by stanbon(75887)

What is the slope of the line containing the points:
(-4, -3), (2, 6), and (6, 12)
----
1st 2 points slope: (6--3)/2--4) = 9/6 = 3/2
-----
last 2 points slope : (12-6)/(6-2) = 6/4 = 3/2
================
Slope = 3/2
===============
Cheers,
Stan H.
============